# Structure-preserving approximation of the non-isothermal Cahn-Hilliard system

### v1

@@ 3 Variational structure
$\langle\partial_{t}\phi,1\rangle=\langle\partial_{t}e,1\rangle=0,\quad\langle\partial_{t}s,1\rangle=\mathcal{D}(\tfrac{\mu}{\vartheta},\tfrac{1}{\vartheta})\geq 0.$
@@
$=\mathcal{D}(\tfrac{\mu}{\vartheta},\tfrac{1}{\vartheta})\geq 0.$
@@
$=0.$
@@ 4 Fully discrete scheme for the non-isothermal Cahn-Hilliard equation
As preparatory step, let us introduce the relevant notation and assumptions for our discretization strategy.
@@
We partition the time interval $[0,T]$ into uniformly divided subintervals with step size $\tau>0$ and introduce $\mathcal{I}_{\tau}:=\{0=t^{0},t^{1}=\tau,\ldots,t^{n_{\tau}}=T\}$, where $n_{\tau}=\tfrac{T}{\tau}$ is the absolute number of time steps. We denote by $\Pi^{1}_{c}(\mathcal{I}_{\tau}),\Pi^{0}(\mathcal{I}_{\tau})$ the spaces of continuous piecewise linear and piecewise constant functions on $\mathcal{I}_{\tau}$. By $g^{n+1},g^{n},g^{*}$ we denote the evaluation of $g\in\Pi^{1}_{c}(\mathcal{I}_{\tau}),\Pi^{0}(\mathcal{I}_{\tau})$ at $t=\{t^{n+1},t^{n},t^{*}\}$, where $t^{*}$ serves as a placeholder for any fixed $t^{*}$ with $t^{*}\in[t^{n},t^{n+1}]$. Further, we introduce the time difference and the discrete time derivative via
@@
For the spacial discretisation we require that the mesh $\mathcal{T}_{h}$ is a geometrically conforming partition of $\Omega$ into simplices of maximal diameter $h$ that can be extended periodically to periodic extensions of $\Omega$. We denote the space of continuous and linear finite element functions over the mesh $\mathcal{T}_{h}$ via
$\mathcal{V}_{h}$ :=$:=\{v\in H^{1}(\Omega)\cap C^{0}(\bar{\Omega}):v|_{K}\in P_{1}(K),\leavevmode\nobreak\ \forall K\in\mathcal{T}_{h}\}.$
We denote the convex-concave splitting of the potential $f$ by
@@
Fully discrete scheme
We propose a fully discrete time-stepping method for the non-isothermal Cahn-Hilliard system (1),(2),(2.3).
@@
$\langle d^{n+1}_{\tau}\phi_{h},\psi_{h}\rangle$ $=\langle\mathbf{M}_{h}^{*}\tfrac{\mu_{h}^{*}}{\vartheta_{h}^{n+1}\cdot\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1}-\tfrac{\mathbf{M}_{h}^{*}}{\vartheta_{h}^{n+1}}\nabla\mu_{h}^{n+1}-\tfrac{\mathbf{C}_{h}^{*}}{\vartheta_{h}^{n+1}\cdot\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1},\nabla\psi_{h}\rangle$ (10)
@@
$\langle d^{n+1}_{\tau}s_{h},\omega_{h}\rangle$ $=\langle\tfrac{K_{h}^{*}}{\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1}+\mathbf{C}_{h}^{*}\nabla\mu_{h}^{n+1}-2\mathbf{C}_{h}^{*}\tfrac{\mu_{h}^{*}}{\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1}-\mathbf{M}_{h}^{*}\mu_{h}^{*}\nabla\mu_{h}^{n+1}$
$\quad+\mathbf{M}_{h}^{*}\tfrac{\mu_{h}^{*}\mu_{h}^{*}}{\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1},\tfrac{\omega_{h}}{\vartheta_{h}^{n+1}\vartheta_{h}^{n+1}\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1}-\tfrac{1}{\vartheta_{h}^{*}\vartheta_{h}^{n+1}}\nabla\omega_{h}\rangle$
$\quad+\langle\tfrac{\mathbf{C}_{h}^{*}}{\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1}+\mathbf{M}_{h}^{*}\nabla\mu_{h}^{n+1}-\mathbf{M}_{h}^{*}\tfrac{\mu_{h}^{*}}{\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1},\tfrac{\omega_{h}}{\vartheta_{h}^{n+1}\vartheta_{h}^{n+1}}\nabla\mu_{h}^{n+1}\rangle$
$\quad+\gamma\langle d^{n+1}_{\tau}\phi_{h}\nabla\phi_{h}^{*},\nabla\omega_{h}\rangle$ (12)
holds for $(\psi_{h},\xi_{h},\omega_{h})\in\mathcal{V}_{h}\times\mathcal{V}_{h}\times\mathcal{V}_{h}$ and all $10\leq n\leq n_{T}-$ 1.
@@ 5 Structure-preserving property
In this section we will show that at least one solution to 5 exists and that the numerical method (10)-(12) preserves the structural properties similar to Theorem 4 at the discrete level.
@@
Every discrete solution $(\phi_{h},\mu_{h},\vartheta_{h})$ of 5 with $\vartheta_{h}>0$ preserves the structural properties, i.e. the conservation of mass, the discrete energy balance and the entropy production
$\langle d^{n+1}_{\tau}\phi_{h},1\rangle$ $=0,\quad\langle d^{n+1}_{\tau}e_{h},1\rangle=\mathcal{D}_{num}^{n+1}\leq 0,\quad\langle d^{n+1}_{\tau}s_{h},1\rangle=\mathcal{D}_{h}^{n+1}\geq 0.$
The discrete entropy production reads
$\mathcal{D}_{h}^{n+1}=\begin{pmatrix}\tfrac{\nabla\vartheta_{h}^{n+1}}{\vartheta_{h}^{n+1}\cdot\vartheta_{h}^{*}}&\tfrac{\mu_{h}^{*}}{\vartheta_{h}^{n+1}\cdot\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1}-\tfrac{\nabla\mu_{h}^{n+1}}{\vartheta_{h}^{n+1}}\end{pmatrix}\begin{pmatrix}\mathbf{K}_{h}^{*}&-\mathbf{C}_{h}^{*}\\ -\mathbf{C}_{h}^{*}&\mathbf{M}_{h}^{*}\end{pmatrix}\begin{pmatrix}\tfrac{\nabla\vartheta_{h}^{n+1}}{\vartheta_{h}^{n+1}\cdot\vartheta_{h}^{*}}\\ \tfrac{\mu_{h}^{*}}{\vartheta_{h}^{n+1}\cdot\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1}-\tfrac{\nabla\mu_{h}^{n+1}}{\vartheta_{h}^{n+1}}\end{pmatrix}$
@@
$\mathcal{D}_{num}^{k}=-\tfrac{\tau}{2}$ $\Big{(}\gamma\langle|\nabla d_{\tau}^{k}\phi_{h}|^{2},\vartheta_{h}^{k}\rangle-\langle\partial_{\vartheta\vartheta}f(\phi_{h}^{k-1},\zeta_{1}),(d_{\tau}^{k}\vartheta_{h})^{2}\rangle$
$+\langle\partial_{\phi\phi}f_{\text{vex}}(\zeta_{2},\vartheta_{h}^{k})-\partial_{\phi\phi}f_{\text{cav}}(\zeta_{3},\vartheta_{h}^{k}),(d_{\tau}^{k}\phi_{h})^{2}\rangle\Big{)}.$
@@
$\quad+\langle\tfrac{\mathbf{C}_{h}^{*}}{\vartheta_{h}^{n+1}\cdot\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1}+\tfrac{\mathbf{M}_{h}^{*}}{\vartheta_{h}^{n+1}}\nabla\mu_{h}^{n+1}-\mathbf{M}_{h}^{*}\tfrac{\mu_{h}^{*}}{\vartheta_{h}^{n+1}\cdot\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1},\nabla\mu_{h}^{n+1}\rangle$
@@
$\quad+\langle\tfrac{\mathbf{C}_{h}^{*}}{\vartheta_{h}^{n+1}\cdot\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1}+\tfrac{\mathbf{M}_{h}^{*}}{\vartheta_{h}^{n+1}}\nabla\mu_{h}^{n+1}-\mathbf{M}_{h}^{*}\tfrac{\mu_{h}^{*}}{\vartheta_{h}^{n+1}\cdot\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1},\nabla\mu_{h}^{n+1}\rangle$
@@
Adding ±$\pm f(\phi_{h}^{n},\vartheta_{h}^{n+1})$ and using (A4) allows us the rewrite as follows
@@ 5.1 Existence of positive solutions
In the next step we will establish existence of discrete solutions with uniformly positive temperature. To this end we need to enhance our assumptions on the driving potential $f$.
@@
[…]
    $\langle e(\phi,\vartheta)-c_{f}s(\phi,\vartheta),1\rangle\leq C_{1}\Longrightarrow\|\nabla\phi\|_{L^{2}}^{2}+\|\vartheta\|_{L^{1}(\Omega)}\leq C_{*}C_{1}.$
@@
    We assume that $|f_{\phi}(\phi_{h}^{n+1},\phi_{h}^{n},\vartheta_{h}^{n+1})|\leq C(|\phi_{h}^{n+1}|^{p_{1}}+|\phi_{h}^{n+1}|^{p_{2}}+|\vartheta_{h}^{n+1}|^{k}+1)$, with $p_{1},p_{2}\leq 6$ and $k\in\mathbb{N}$.
@@
To prove the existence result it suffices to consider a single time step. For simplicity we consider $g^{*}=g^{n}$ for all $g\in\{\phi_{h},\mu_{h}\vartheta_{h}\}$ in 5, but the whole proof works for other choices of $g^{*}$. After choosing the Lagrange basis in $\mathcal{V}_{h}$ we can rewrite 5 as a finite-dimensional nonlinear problem $\mathbf{g}(\mathbf{x})=\mathbf{0}$, where $\mathbf{x}=(\boldsymbol{\phi},\boldsymbol{\mu},\boldsymbol{\vartheta})\in\mathbb{R}^{3N}$ is a vector containing the nodal values of $x_{h}:=(\phi^{n+1}_{h},\mu_{h}^{n+1},\vartheta_{h}^{n+1})$. For simplicity we will use the abbreviation $(\phi_{h},\mu_{h},\vartheta_{h})\equiv(\phi_{h}^{n+1},\mu_{h}^{n+1},\vartheta_{h}^{n+1})$. To apply Theorem 1 we consider
$V$ :=$:=\{(\boldsymbol{\phi},\boldsymbol{\mu},\boldsymbol{\vartheta})\in\mathbb{R}^{3N}\text{ such that }\boldsymbol{\vartheta}>0\},$ (15)
$W$ :=$:=\{(\boldsymbol{\phi},\boldsymbol{\mu},\boldsymbol{\vartheta})\in\mathbb{R}^{3N}\text{ such that }\|\boldsymbol{\phi}\|+\|\boldsymbol{\mu}\|\leq C_{\phi,\mu},\varepsilon<\boldsymbol{\vartheta}<C_{\vartheta},\}.$ (16)
@@
In the next step we will define $\mathbf{G}(\mathbf{x},\alpha)=\langle T_{\alpha}x_{h},y_{h}\rangle$ with $\langle T_{\alpha}x_{h},y_{h}\rangle$ given by the variational problem
@@
$\frac{1}{\tau}\langle\phi_{h}-\phi_{h}^{n},\psi_{h}\rangle-\alpha\langle\tfrac{\mathbf{M}_{h}^{n}}{\vartheta_{h}^{n+1}\cdot\vartheta_{h}^{n}}\mu_{h}^{n}\nabla\vartheta_{h}-\tfrac{\mathbf{M}_{h}^{n}}{\vartheta_{h}}\nabla\mu_{h}-\tfrac{\mathbf{C}_{h}^{n}}{\vartheta_{h}\cdot\vartheta_{h}^{n}}\nabla\vartheta_{h},\nabla\psi_{h}\rangle=0$ (17)
$\langle\mu_{h},\xi_{h}\rangle-\gamma\langle\nabla\phi_{h},\vartheta_{h}\nabla\xi_{h}\rangle-\gamma\langle\nabla\phi_{h}^{n},\xi_{h}\nabla\vartheta_{h}\rangle-\langle\partial_{\phi}f_{\phi}(\phi_{h},\phi_{h}^{n},\vartheta_{h}),\xi_{h}\rangle=0$ (18)
@@
$\frac{1}{\tau}\langle s(\phi_{h},\vartheta_{h})-s(\phi_{h}^{n},\vartheta_{h}^{n}),\omega_{h}\rangle-\alpha\langle\tfrac{\mathbf{K}_{h}^{n}}{\vartheta_{h}\cdot\vartheta_{h}^{n}}\nabla\vartheta_{h}+\tfrac{\mathbf{C}_{h}^{n}}{\vartheta_{h}}\nabla\mu_{h}-2\tfrac{\mathbf{C}_{h}^{n}}{\vartheta_{h}\cdot\vartheta_{h}^{n}}\mu_{h}^{n}\nabla\vartheta_{h}$
$\quad-\mathbf{M}_{h}^{n}\tfrac{\mu_{h}^{n}}{\vartheta_{h}}\nabla\mu_{h}+\mathbf{M}_{h}^{n}\tfrac{\left(\mu_{h}^{n}\right)^{2}}{\vartheta_{h}\cdot\vartheta_{h}^{n}}\nabla\vartheta_{h},\tfrac{\omega_{h}}{\vartheta_{h}\cdot\vartheta_{h}^{n}}\nabla\vartheta_{h}-\tfrac{1}{\vartheta_{h}^{n}}\nabla\omega_{h}\rangle$
@@
$\quad-\alpha\langle\tfrac{\mathbf{C}_{h}^{n}}{\vartheta_{h}\cdot\vartheta_{h}^{n}}\nabla\vartheta_{h}+\tfrac{\mathbf{M}_{h}^{n}}{\vartheta_{h}}\nabla\mu_{h}-\tfrac{\mathbf{M}_{h}^{n}}{\vartheta_{h}\cdot\vartheta_{h}^{n}}\mu_{h}^{n}\nabla\vartheta_{h},\tfrac{\omega_{h}}{\vartheta_{h}}\nabla\mu_{h}\rangle$
$\quad+\frac{1}{\tau}\gamma\langle\nabla\phi_{h}^{n}(\phi_{h}-\phi_{h}^{n}),\nabla\omega_{h}\rangle=0$ (19)

### v2

@@ 3 Variational structure
$\langle\partial_{t}\phi,1\rangle=\langle\partial_{t}e,1\rangle=0,\quad\langle\partial_{t}s,1\rangle=\mathcal{D}\left(\tfrac{\mu}{\vartheta},\tfrac{1}{\vartheta}\right)\geq 0.$
@@
$=\mathcal{D}\left(\tfrac{\mu}{\vartheta},\tfrac{1}{\vartheta}\right)\geq 0.$
@@
$\quad+\langle\tfrac{\mathbf{C}}{\vartheta^{2}}\nabla\vartheta+\tfrac{\mathbf{M}}{\vartheta}\nabla\mu-\mathbf{M}\tfrac{\mu}{\vartheta^{2}}\nabla\vartheta,\nabla\mu\rangle=0.$
@@ 4 Fully discrete scheme for the non-isothermal Cahn-Hilliard equation
As preparatory step, we introduce the notation and assumptions that will be used in our discretisation strategy.
@@
We consider a uniform partition of the time interval $[0,T]$ with time step size $\tau>0$ and define the discrete time grid $\mathcal{I}_{\tau}:=\{0=t^{0},t^{1}=\tau,\ldots,t^{n_{\tau}}=T\}$, where $n_{\tau}=\tfrac{T}{\tau}$ denotes the absolute number of time steps. We introduce $\Pi^{1}_{c}(\mathcal{I}_{\tau};X),\Pi^{0}(\mathcal{I}_{\tau};X)$ as the spaces of continuous piecewise linear and piecewise constant functions on $\mathcal{I}_{\tau}$ with values in a suitable Hilbert space $X$, respectively. For a function $g\in\Pi^{1}_{c}(\mathcal{I}_{\tau};X),\Pi^{0}(\mathcal{I}_{\tau};X)$, we use the notation $g^{n+1},g^{n},g^{*}$ to refer to its evaluation at $t^{n+1},t^{n}$, and at some intermediate time $t^{*}\in[t^{n},t^{n+1}]$, respectively. The time difference and the backward finite difference quotient (discrete time derivative) are defined by
@@
For the spatial discretization, we consider a geometrically conforming triangulation $\mathcal{T}_{h}$ of the domain $\Omega$ into simplices with maximal diameter $h.$ The mesh is assumed to be periodically extendable to match the periodic structure of $\Omega$. We define the space of continuous, piecewise linear finite element functions over $\mathcal{T}_{h}$ by
$\mathcal{V}_{h}$ $:=\{v\in H^{1}(\Omega)\cap C^{0}(\bar{\Omega}):v|_{K}\in P_{1}(K),\penalty 10000\ \forall K\in\mathcal{T}_{h}\}.$
For $\phi_{h},\vartheta_{h}\in\Pi^{1}_{c}(\mathcal{I}_{\tau};\mathcal{V}_{h})$ and $n\in\{0,...,n_{\tau}-1\}$ we adopt a convex-concave splitting with respect to the phase variable $\phi$ of the potential $f$, and define the discrete approximation of its derivative as
@@
For the components of the Onsager matrix $\mathbf{X}\in\{\mathbf{M},\mathbf{K},\mathbf{C}\}$ we denote by $\mathbf{X}_{h}^{*}$ the evaluation at $(\phi_{h}^{*},\nabla\phi_{h}^{*},\vartheta_{h}^{*})$ for some intermediate time $t^{*}\in[t^{n},t^{n+1}]$, i.e. $\mathbf{X}_{h}^{*}=\mathbf{X}(\phi_{h}^{*},\nabla\phi_{h}^{*},\vartheta_{h}^{*})$.
We now formulate a fully discrete time-stepping method for the non-isothermal Cahn–Hilliard system (1), (2), and (2.3).
@@
$\langle d^{n+1}_{\tau}\phi_{h},\psi_{h}\rangle$ $=\langle\tfrac{\mu_{h}^{*}}{\vartheta_{h}^{*}}\mathbf{M}_{h}^{*}\nabla\vartheta_{h}^{n+1}-\mathbf{M}_{h}^{*}\nabla\mu_{h}^{n+1}-\tfrac{1}{\vartheta_{h}^{*}}\mathbf{C}_{h}^{*}\nabla\vartheta_{h}^{n+1},\tfrac{1}{\vartheta_{h}^{n+1}}\nabla\psi_{h}\rangle$ (14)
@@
$\langle d^{n+1}_{\tau}s_{h},\omega_{h}\rangle$ $=\langle\tfrac{1}{\vartheta_{h}^{*}}\mathbf{K}_{h}^{*}\nabla\vartheta_{h}^{n+1}+\mathbf{C}_{h}^{*}\nabla\mu_{h}^{n+1}-2\tfrac{\mu_{h}^{*}}{\vartheta_{h}^{*}}\mathbf{C}_{h}^{*}\nabla\vartheta_{h}^{n+1}-\mu_{h}^{*}\mathbf{M}_{h}^{*}\nabla\mu_{h}^{n+1}$
$\quad+\tfrac{(\mu_{h}^{*})^{2}}{\vartheta_{h}^{*}}\mathbf{M}_{h}^{*}\nabla\vartheta_{h}^{n+1},\tfrac{\omega_{h}}{(\vartheta_{h}^{n+1})^{2}\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1}-\tfrac{1}{\vartheta_{h}^{*}\vartheta_{h}^{n+1}}\nabla\omega_{h}\rangle$
$\quad+\langle\tfrac{1}{\vartheta_{h}^{*}}\mathbf{C}_{h}^{*}\nabla\vartheta_{h}^{n+1}+\mathbf{M}_{h}^{*}\nabla\mu_{h}^{n+1}-\tfrac{\mu_{h}^{*}}{\vartheta_{h}^{*}}\mathbf{M}_{h}^{*}\nabla\vartheta_{h}^{n+1},\tfrac{\omega_{h}}{(\vartheta_{h}^{n+1})^{2}}\nabla\mu_{h}^{n+1}\rangle$
$\quad+\gamma\langle\nabla\phi_{h}^{*},d^{n+1}_{\tau}\phi_{h}\nabla\omega_{h}\rangle$ (16)
holds for $(\psi_{h},\xi_{h},\omega_{h})\in\mathcal{V}_{h}\times\mathcal{V}_{h}\times\mathcal{V}_{h}$ and all $0\leq n\leq n_{T}-1$.
@@ 5 Structure-preserving property
In this section, we establish the existence of at least one solution to 5, and we demonstrate that the numerical method (14)–(16) preserves key structural properties at the discrete level—analogous to those described in Theorem 4. In particular, we verify discrete analogues of mass conservation, energy dissipation, and entropy production.
@@
Let $(\phi_{h},\mu_{h},\vartheta_{h})$ be any discrete solution of 5 with $\vartheta_{h}>0.$ Then the scheme preserves the following structural properties:
$\langle d^{n+1}_{\tau}\phi_{h},1\rangle$ $=0,\qquad$ (discrete mass conservation)
$\langle d^{n+1}_{\tau}e_{h},1\rangle$ $=\mathcal{D}^{n+1}_{\mathrm{num}}\leq 0,\qquad$ (discrete energy dissipation)
$\langle d^{n+1}_{\tau}s_{h},1\rangle$ $=\mathcal{D}_{h}^{n+1}\geq 0,\qquad$ (discrete entropy production)
where $e_{h}$ and $s_{h}$ denote the discrete internal energy and entropy, respectively. The discrete entropy production reads
$\mathcal{D}_{h}^{n+1}=\int_{\Omega}\begin{pmatrix}\tfrac{\nabla\vartheta_{h}^{n+1}}{\vartheta_{h}^{n+1}\cdot\vartheta_{h}^{*}}&\tfrac{\mu_{h}^{*}}{\vartheta_{h}^{n+1}\cdot\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1}-\tfrac{\nabla\mu_{h}^{n+1}}{\vartheta_{h}^{n+1}}\end{pmatrix}\begin{pmatrix}\mathbf{K}_{h}^{*}&-\mathbf{C}_{h}^{*}\\ -\mathbf{C}_{h}^{*}&\mathbf{M}_{h}^{*}\end{pmatrix}\begin{pmatrix}\tfrac{\nabla\vartheta_{h}^{n+1}}{\vartheta_{h}^{n+1}\cdot\vartheta_{h}^{*}}\\ \tfrac{\mu_{h}^{*}}{\vartheta_{h}^{n+1}\cdot\vartheta_{h}^{*}}\nabla\vartheta_{h}^{n+1}-\tfrac{\nabla\mu_{h}^{n+1}}{\vartheta_{h}^{n+1}}\end{pmatrix}$
@@
$\mathcal{D}_{num}^{n+1}=-\tfrac{\tau}{2}$ $\Big(\gamma\langle|\nabla d_{\tau}^{n+1}\phi_{h}|^{2},\vartheta_{h}^{n+1}\rangle-\langle\partial_{\vartheta\vartheta}f(\phi_{h}^{n},\zeta_{1}),(d_{\tau}^{n+1}\vartheta_{h})^{2}\rangle$
$+\langle\partial_{\phi\phi}f_{\text{vex}}(\zeta_{2},\vartheta_{h}^{{n+1}})-\partial_{\phi\phi}f_{\text{cav}}(\zeta_{3},\vartheta_{h}^{{n+1}}),(d_{\tau}^{n+1}\phi_{h})^{2}\rangle\Big).$
@@
$\quad+\langle\tfrac{1}{\vartheta_{h}^{*}}\mathbf{C}_{h}^{*}\nabla\vartheta_{h}^{n+1}+\mathbf{M}_{h}^{*}\nabla\mu_{h}^{n+1}-\tfrac{\mu_{h}^{*}}{\vartheta_{h}^{*}}\mathbf{M}_{h}^{*}\nabla\vartheta_{h}^{n+1},\tfrac{1}{\vartheta_{h}^{n+1}}\nabla\mu_{h}^{n+1}\rangle$
@@
$\quad+\langle\tfrac{1}{\vartheta_{h}^{*}}\mathbf{C}_{h}^{*}\nabla\vartheta_{h}^{n+1}+\mathbf{M}_{h}^{*}\nabla\mu_{h}^{n+1}-\tfrac{\mu_{h}^{*}}{\vartheta_{h}^{*}}\mathbf{M}_{h}^{*}\nabla\vartheta_{h}^{n+1},\tfrac{1}{\vartheta_{h}^{n+1}}\nabla\mu_{h}^{n+1}\rangle$
@@
Adding $\pm f(\phi_{h}^{n},\vartheta_{h}^{n+1})$ and using (A4) allows us the rewrite as follows
@@ 5.1 Existence of positive solutions
In the next step, we will establish existence of discrete solutions with uniformly positive temperature. To this end we need to restrict the assumptions on the potential $f$.
@@
    We extend (A4) and assume that $f$ is of the form
    $f(\phi,\vartheta)=-b\vartheta\log(\vartheta)+\vartheta f_{1}(\phi)+f_{2}(\phi)$
    with $f_{2}(\phi)\geq-c$ for some constants $b,c>0$ independent of $\phi$ and $\vartheta$. Furthermore, we assume that there exists another constant $c_{f}>0$ independent of $\phi$ such that
    $f_{2}(\phi)+c_{f}f_{1}(\phi)\geq-c.$
    By construction $F(\phi,\cdot):\mathbb{R}_{+}\to\mathbb{R}$ is strictly concave.
@@
    We assume that the convex-concave decomposition of $f$ fulfills
    $|f_{\phi}(\phi_{h}^{n+1},\phi_{h}^{n},\vartheta_{h}^{n+1})|\leq C_{f}(|\vartheta_{h}^{n+1}|+1)(|\phi_{h}^{n+1}|^{6}+|\phi_{h}^{n}|^{6}+1)$
    for some constant $C_{f}>0$.
@@
Remark 7.
Assumption (A5) is satisfied for
$\quad-b(\vartheta\log(\tfrac{\vartheta}{\vartheta_{c}})+\vartheta-\vartheta_{c})$
for $c_{f}=\vartheta_{c},b=b,c=\min\{\vartheta_{c}\log(\vartheta_{c}),0\}$ given that $a,b,d,\vartheta_{c}>0$, see also [3] for a similar assumption.
@@
To prove the existence result it suffices to consider a single time step $0\leq n\rightarrow n+1\leq n_{\tau}$. For simplicity we consider $g^{*}=g^{n}+1$ and $\mathbf{X}_{h}^{*}=\mathbf{X}_{h}^{n}+1$ for all $g\in\{\phi_{h},\mu_{h}\vartheta_{h}\}$ and $\mathbf{X}\in\{\mathbf{M},\mathbf{K},\mathbf{C}\}$ in 5, but the whole proof works for other choices of $g^{*}$ and $\mathbf{X}_{h}^{*}$. After choosing the Lagrange basis in $\mathcal{V}_{h}$ we can rewrite 5 as a finite-dimensional nonlinear problem $\mathbf{g}(\mathbf{x})=\mathbf{0}$, i.e. $\mathbf{b}=0$, where $\mathbf{x}=(\boldsymbol{\phi},\boldsymbol{\mu},\boldsymbol{\vartheta})\in\mathbb{R}^{3N}$ is a vector containing the nodal values of $(\phi^{n+1}_{h},\mu_{h}^{n+1},\vartheta_{h}^{n+1})$. For simplicity we will use the abbreviation $(\phi_{h},\mu_{h},\vartheta_{h})\equiv(\phi_{h}^{n+1},\mu_{h}^{n+1},\vartheta_{h}^{n+1})$ as well as $\mathbf{X}_{h}\equiv\mathbf{X}_{h}^{n+1}$. To apply Theorem 1 we consider
$V$ $:=\{(\boldsymbol{\phi},\boldsymbol{\mu},\boldsymbol{\vartheta})\in\mathbb{R}^{3N}\text{ such that }\boldsymbol{\vartheta}>0\},$ (19)
$W$ $:=\{(\boldsymbol{\phi},\boldsymbol{\mu},\boldsymbol{\vartheta})\in\mathbb{R}^{3N}\text{ such that }\|\boldsymbol{\phi}\|+\|\boldsymbol{\mu}\|<C_{\phi,\mu},\varepsilon<\boldsymbol{\vartheta}<C_{\vartheta},\}.$ (20)
@@
First, we define $\mathbf{G}(\mathbf{x},\alpha):=\langle T_{\alpha}x_{h},y_{h}\rangle=\mathbf{0}$ for $x_{h}=(\phi_{h},\mu_{h},\vartheta_{h})\in V$ and all basis functions $y_{h}=(\psi_{h},\xi_{h},\omega_{h})\in\mathcal{V}_{h}\times\mathcal{V}_{h}\times\mathcal{V}_{h}$ with $\langle T_{\alpha}x_{h},y_{h}\rangle=0$ given by the variational problem
@@
$\frac{1}{\tau}\langle\phi_{h}-\phi_{h}^{n},\psi_{h}\rangle-\alpha\langle\tfrac{\mathbf{M}_{h}}{\vartheta_{h}^{2}}\mu_{h}\nabla\vartheta_{h}-\tfrac{\mathbf{M}_{h}}{\vartheta_{h}}\nabla\mu_{h}-\tfrac{\mathbf{C}_{h}}{\vartheta_{h}^{2}}\nabla\vartheta_{h},\nabla\psi_{h}\rangle=0$ (21)
$\langle\mu_{h},\xi_{h}\rangle-\gamma\langle\nabla\phi_{h},\vartheta_{h}\nabla\xi_{h}\rangle-\gamma\langle\nabla\phi_{h},\xi_{h}\nabla\vartheta_{h}\rangle-\langle f_{\phi}(\phi_{h},\phi_{h}^{n},\vartheta_{h}),\xi_{h}\rangle=0$ (22)
@@
$\frac{1}{\tau}\langle s(\phi_{h},\vartheta_{h})-s(\phi_{h}^{n},\vartheta_{h}^{n}),\omega_{h}\rangle-\alpha\langle\tfrac{\mathbf{K}_{h}}{\vartheta_{h}^{2}}\nabla\vartheta_{h}+\tfrac{\mathbf{C}_{h}}{\vartheta_{h}}\nabla\mu_{h}-2\tfrac{\mathbf{C}_{h}}{\vartheta_{h}^{2}}\mu_{h}\nabla\vartheta_{h}$
$\quad-\mathbf{M}_{h}\tfrac{\mu_{h}}{\vartheta_{h}}\nabla\mu_{h}+\mathbf{M}_{h}\tfrac{\mu_{h}}{\vartheta_{h}}^{2}\nabla\vartheta_{h},\tfrac{\omega_{h}}{\vartheta_{h}^{2}}\nabla\vartheta_{h}-\tfrac{1}{\vartheta_{h}}\nabla\omega_{h}\rangle$
@@
$\quad-\alpha\langle\tfrac{\mathbf{C}_{h}}{\vartheta_{h}^{2}}\nabla\vartheta_{h}+\tfrac{\mathbf{M}_{h}}{\vartheta_{h}}\nabla\mu_{h}-\tfrac{\mathbf{M}_{h}}{\vartheta_{h}^{2}}\mu_{h}\nabla\vartheta_{h},\tfrac{\omega_{h}}{\vartheta_{h}}\nabla\mu_{h}\rangle$
$\quad+\frac{1}{\tau}\gamma\langle\nabla\phi_{h}(\phi_{h}-\phi_{h}^{n}),\nabla\omega_{h}\rangle=0$ (23)